A cannon on top of a 12 meter castle wall fires a cannonball horizonally at an initial speed () I the cannonball hits the perfectly flat ground 93 meters away from the castle wall, what was the initial speed of the cannonball? Value Units
Answer #1

considering motion along vertical

h = 0.5* gt^2

12 = 0.5* 9.8* t^2

t = 1.565 s

Considering motion along horizontal

X = vo * t

93 = vo * 1.565

vo = 59.428 m/s

=========

do comment in case any doubt, will reply for sure . Goodluck